563blackghost (563blackghost):

The graph \(\large{g(x)}\) is the graph of \(\large{f(x)=x^{2}}\) shifted 4 units left, vertically stretched by a factor of 3, then shifted 5 units up. What is the function rule for \(\large{g(x)}\)? I know how to apply the translation but I need help applying the stretch.

563blackghost (563blackghost):

Now I currently have \(\Large{g(x+4)+5}\) But i don't know on how to apply the stretch of 3.

OpenStudy (3mar):

May I help, sister?

563blackghost (563blackghost):

Sure :)

OpenStudy (3mar):

Thanks

563blackghost (563blackghost):

Is what I have currently correct? @3mar

OpenStudy (3mar):

Yes, without the stretch.

563blackghost (563blackghost):

Great :)

563blackghost (563blackghost):

Would with the stretch applied be like this? \(\huge{y=5g(x+4)+5}\) According to my textbook it says that... (sorry my internet is not so good at the moment)

OpenStudy (3mar):

Never mind. I think these are the steps.

OpenStudy (3mar):

Let's test your result.

563blackghost (563blackghost):

Oooo I forgot the square :(

563blackghost (563blackghost):

\(\huge{y=5g(x+4)^{2}+5}\)

563blackghost (563blackghost):

I mean't 3g sorry....

563blackghost (563blackghost):

Wait g will not be part of the equation? it would be in place of y?

OpenStudy (3mar):

But I think another way. can you discuss it with me?

563blackghost (563blackghost):

Yes :)

OpenStudy (3mar):

Ok. Let's.

563blackghost (563blackghost):

So we would need to find g(x) from \(\large{f(x)=x^{2}}\) we would first apply the translation of 4 to the left and 5 up...in which the equation then turns to... \(\Large{g(x)=(x+4)^{2}+5}\) Correct?

OpenStudy (3mar):

Parent function: \[f(x)=x^2\] 4 unit shift left : \[f(x)=(x+4)^2\] and \[f(x)=g(x+4)\] 3 factors vertical stretch: \[f(x)=3(x+4)^2\] and \[f(x)=g(\sqrt{3}( x+4))\] finally, 5 units shift up: \[f(x)=3(x+4)^2+5\] and \[f(x)=g(\sqrt{3}(x+4))+5\] Hope I am right!

OpenStudy (3mar):

Take care: "we would first apply the translation of 4 to the left and 5 up" Order of step is important and it will make a difference for other problems if it does not make her.

563blackghost (563blackghost):

I don't quite understand as to how 3 is rooted.

563blackghost (563blackghost):

Ooo I see my mistake I mean't it as `what i know first` maybe I shouldn't have worded it like that. Thank you for pointing that out @3mar :)

563blackghost (563blackghost):

I seem to understand as how the function is written by studying it quite a bit. Thank you so much @3mar :)

OpenStudy (3mar):

Wait, sister! You said that:" y=5g(x+4)+5 According to my textbook it says that..."

OpenStudy (3mar):

I just wonder how we substitute with (x+4) and then *5?

563blackghost (563blackghost):

No, that was before but I now understand my mistake @3mar thanks to you :)

OpenStudy (3mar):

Every thing is ok?

563blackghost (563blackghost):

Yes @3mar Thank you so much :)

OpenStudy (3mar):

But I might accept the form: \[g(x)=3f(x+4)+5\]? ? It is quiet right. But I don't want to get you confused If you are, just make sure from your teacher. and kindly let me know which one is more accurate.

563blackghost (563blackghost):

I will @3mar :) I'll make sure of it.

OpenStudy (3mar):

If there is more, I won't be late!

563blackghost (563blackghost):

No, that is all for today. Thankfully :) Once again, Thank you @3mar

OpenStudy (3mar):

You are welcome, sister! Be well all the time.